OpenStudy (anonymous):

Find the number of real solutions of \(\Large \frac{x^4-20x^3+150x^2-500x+625}{x-5}=0\)

OpenStudy (dan815):

ok hmmm lets see!!

OpenStudy (dan815):

see if x-5 is factorable first

OpenStudy (dan815):

because x=5 might infact be defined

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

\[\large{x^4-20x^3+150x^2-500x+625}\] \[\large{ = x^3(x-5) - 15x^2(x-5) + 75x(x-5)}\] \[\large{ - 125(x-5)}\] \[\large{= (x-5)(x^3-15x^2+75x-125)}\]

OpenStudy (dan815):

oops x=5 is good :)

OpenStudy (dan815):

infact its good again for the x^3 func too

OpenStudy (vishweshshrimali5):

Actually it is very very good. Its finally (x-5)^4

OpenStudy (dan815):

right so answer is (x-5)^3

OpenStudy (anonymous):

5 is incorrect

OpenStudy (vishweshshrimali5):

\[(x-5)(x^3-15x^2+75x-125)\] \[= (x-5)[x^2(x-5) - 10x(x-5) + 25(x-5)]\] \[= (x-5)^2[x^2-10x+25]\] \[= (x-5)^2(x-5)^2\] \[=(x-5)^4\]

OpenStudy (vishweshshrimali5):

No real solution exists

OpenStudy (vishweshshrimali5):

Because we can only cancel out (x-5) in denominator and numerator if \(x\ne 5\)/

OpenStudy (dan815):

oh makes sense

OpenStudy (vishweshshrimali5):

Now our equation would become: \[(x-5)^3 = 0 ;~~~~x\ne 5\] \[\implies x = 5 ;~~~~~~~x\ne 5\] Thus, no solution

OpenStudy (dan815):

then what happens at x=5?? do we say a hole ?

OpenStudy (vishweshshrimali5):

Yep

OpenStudy (vishweshshrimali5):

We get a 0/0 form

OpenStudy (vishweshshrimali5):

Function is non continuous at x = 5

OpenStudy (dan815):

quessioon tho

OpenStudy (anonymous):

So, here are no real solutions and the answer is \(\Large 0\) :)

OpenStudy (vishweshshrimali5):

Yep :)

OpenStudy (dan815):

supposed u have x^2=0 now i can multiply x/x * x^2=0 does this mean.., for x=0 its undefined now

OpenStudy (vishweshshrimali5):

Actually x/x would make your solution wrong

OpenStudy (dan815):

soo why cant this just be a (x-5)^3 cube shifted 5 to the right

OpenStudy (dan815):

how come

OpenStudy (dan815):

is it because u cant say some number over some number =1 since 0/0 and inf/inf or 0/inf is undefined

OpenStudy (vishweshshrimali5):

See, using x/x is possible only when \(x \ne 0\)

OpenStudy (dan815):

hmm ok

OpenStudy (dan815):

goood to knoww

OpenStudy (vishweshshrimali5):

For example taking square roots on LHS and RHS is possible only when you are sure that none of them is negative or we would gets stuck :)

OpenStudy (vishweshshrimali5):

(x^2-6) = -15 If I take square roots on both sides in this step, I am making a wrong step and I would get no solution

OpenStudy (vishweshshrimali5):

But if I solve it a little bit : x^2 = 9 At this step taking square root is correct and permissible. I would get 2 solutions 3, -3

OpenStudy (vishweshshrimali5):

So, from one step we were getting no solutions but from the other we are getting 2 solutions. Only because we were using a wrong operation :)